The area of a rectangle is 45 inches. One dimension is 5 inches. What is the perimeter?
Answer by mananth(16946)

The area of a rectangle is 45 inches. One dimension is 5 inches. What is the perimeter?
Area = L*W
45 = L*5
L=45/5
L=9 inches
Perimeter = 2L + 2W length = 9 and width =5
= 2*9 +2*5
= 18+10
= 28 inches.